# Towards Controllable Speech Synthesis in the Era of Large Language Models: A Survey

### v1

@@ Abstract
Text-to-speech (TTS), also known as speech synthesis, is a prominent research area that aims to generate natural-sounding human speech from text.  Recently, with the increasing industrial demand, TTS technologies have evolved beyond synthesizing human-like speech to enabling controllable speech generation. This includes fine-grained control over various attributes of synthesized speech such as emotion, prosody, timbre, and duration. Besides, advancements in deep learning, such as diffusion and large language models, have significantly enhanced controllable TTS over the past several years. In this paper, we conduct a comprehensive survey of controllable TTS, covering approaches ranging from basic control techniques to methods utilizing natural language prompts, aiming to provide a clear understanding of the current state of research. We examine the general controllable TTS pipeline, challenges, model architectures, and control strategies, offering a comprehensive and clear taxonomy of existing methods. Additionally, we provide a detailed summary of datasets and evaluation metrics and shed some light on the applications and future directions of controllable TTS. To the best of our knowledge, this survey paper provides the first comprehensive review of emerging controllable TTS methods, which can serve as a beneficial resource for both academic researchers and industry practitioners.
@@ I Introduction
Deep learning [12] has made great progress in the past decade due to exponentially growing computational resources like GPUs [13], leading to the explosion of numerous great works on TTS [14, 15, 16, 17]. These methods can synthesize human speech with better quality [14] and can achieve fine-grained control of the generated voice [18, 19, 20, 21, 22]. Besides, some recent works synthesize speech given multi-modal input, such as face images [23, 24], cartoons [7], and videos [25]. Moreover, with the fast development of open-source LLMs [11, 26, 27, 28, 29], some researchers propose to synthesize fine-grained controllable speech with natural language description [30, 31, 32], coining a new way to generate custom speech voices. Meanwhile, powering LLMs with speech synthesis has also been a hot topic in the last few years [33, 34, 35]. In recent years, a wide range of TTS methods has emerged, making it essential for researchers to gain a comprehensive understanding of current research trends, particularly in controllable TTS, to identify promising future directions in this rapidly evolving field. Consequently, there is a pressing need for an up-to-date survey of TTS techniques. While several existing surveys address parametric-based approaches [36, 37, 38, 39, 40, 41] and deep learning-based TTS [42, 43, 44, 45, 46, 47, 48], they largely overlook the controllability of TTS. Additionally, these surveys do not cover the advancements in recent years, such as natural language description-based TTS methods.
This paper provides a comprehensive and in-depth survey of existing and emerging TTS technologies, with a particular focus on controllable TTS methods. Fig. 1 demonstrates the development of controllable TTS methods in recent years, showing their backbones, feature representations, and control abilities. The remainder of this section begins with a brief comparison between this survey and previous ones, followed by an overview of the history of controllable TTS technologies, ranging from early milestones to state-of-the-art advancements. Finally, we introduce the taxonomy and organization of this paper.
Figure 1: A summary of representative controllable TTS methods in recent years and their model architectures, feature representations, and control abilities. There are additional network structures, such as VAE-based models, that are not included in this figure. For more details, refer to Tables IV and III.
@@ I-A Comparison with Existing Surveys
Different scope. Klatt et al. [36] provided the first comprehensive survey on formant, concatenative, and articulatory TTS methods, with a strong emphasis on text analysis. In the early 2010s, Tabet et al. [49] and King et al. [40] explored rule-based, concatenative, and HMM-based techniques. Later, the advent of deep learning catalyzed the emergence of numerous neural-based TTS methods. Therefore, Ning et al. [43] and Tan et al. [42] have conducted extensive surveys on neural-based acoustic models and vocoders, while Zhang et al. [50] presented the first review of diffusion model-based TTS techniques. However, these studies offer limited discussion on the controllability of TTS systems. To address this gap, we present the first comprehensive survey of TTS methods through the lens of controllability, providing an in-depth analysis of model architectures and strategies for controlling synthesized speech.
Close to current demand. With the rapid development of hardware (i.e., GPUs) and artificial intelligence techniques (i.e., transformers, LLMs, diffusion models) in the last few years, the demand for controllable TTS is becoming increasingly urgent due to its broad applications in industries such as filmmaking, gaming, robots, and personal assistants. Despite this growing need, existing surveys pay little attention to control methods in TTS technologies. To bridge this gap, we propose a systematic analysis of current controllable TTS methods and the associated challenges, offering a comprehensive understanding of the research state in this field.
@@
Table I summarizes representative surveys and this paper in terms of main focus and publication year.
@@
Tan et al. [42] Neural-based, efficient, and expressive TTS 2021
@@
Ours Controllable TTS and evaluation 2024
@@ I-B The History of Controllable TTS
Controllable TTS aims to control various aspects of synthesized speech, such as pitch, energy, speed/duration, prosody, timbre, emotion, gender, or high-level styles. This subsection briefly reviews the history of controllable TTS ranging from early approaches to the state-of-arts in recent years.
@@
Early approaches. Before the prevalence of deep neural networks (DNNs), controllable TTS technologies were built primarily on rule-based, concatenative, and statistical methods. These approaches enable some degree of customization and control, though they were constrained by the limitations of the underlying models and available computational resources. 1) Rule-based TTS systems [51, 52, 53, 54], such as formant synthesis, were among the earliest methods for speech generation. These systems use manually crafted rules to simulate the speech generation process by controlling acoustic parameters such as pitch, duration, and formant frequencies, allowing explicit manipulation of prosody and phonetic details through rule adjustments. 2) Concatenative TTS [55, 56, 57, 58], which dominated the field in the late 1990s and early 2000s, synthesize speech by concatenating pre-recorded speech segments, such as phonemes or diphones, stored in a large database [59]. These methods can modify the prosody by manipulating the pitch, duration, and amplitude of speech segments during concatenation. They also allow limited voice customization by selecting speech units from different speakers. 3) Parametric methods, particularly HMM-based TTS [60, 61, 62, 63, 64, 65], gained prominence in the late 2000s. These systems model the relationships between linguistic features and acoustic parameters, providing more flexibility in controlling prosody, pitch, speaking rate, and timbre by adjusting statistical parameters. Some HMM-based systems also supported speaker adaptation [66, 67] and voice conversion [68, 69], enabling voice cloning to some extent. Besides, emotion can also be limitedly controlled by some of these methods [70, 71, 60, 72]. In addition, they required less storage compared to concatenative TTS and allowed smoother transitions between speech units.
@@
Neural-based synthesis. Neural-based TTS technologies emerged with the advent of deep learning, significantly advancing the field by enabling more flexible, natural, and expressive speech synthesis. Unlike traditional methods, neural-based TTS leverages DNNs to model complex relationships between input text and speech, facilitating nuanced control over various speech characteristics. Early neural TTS systems, such as WaveNet [73] and Tacotron [74] laid the groundwork for controllability. 1) Controlling prosody features like rhythm and intonation is vital for generating expressive and contextually appropriate speech. Neural-based TTS models achieve prosody control through explicit conditioning or learned latent representations [75, 15, 76, 77, 78]. 2) Speaker control has also gained significant improvement in neural-based TTS through speaker embeddings or adaptation techniques [79, 80, 81, 82]. 3) Besides, emotionally controllable TTS [83, 31, 20, 22, 32] has become a hot topic due to the strong modeling capability of DNNs, enabling the synthesis of speech with specific emotional tones such as happiness, sadness, anger, or neutrality. These systems go beyond producing intelligible and natural-sounding speech, focusing on generating expressive output that aligns with the intended emotional context. 4) Neural-based TTS can also manipulate timbre (vocal quality) [78, 84, 85, 14, 86, 87] and style (speech mannerisms) [88, 89, 90], allowing for creative and personalized applications. These techniques lead to one of the most popular research topics, i.e., zero-shot TTS (particularly voice cloning) [82, 78, 91, 92]. 5) Fine-grained content and linguistic control also become more powerful [93, 94, 95, 96]. These methods can emphasize or de-emphasize specific words or adjust the pronunciation of phonemes through speech editing or generation techniques.
@@
Neural-based TTS technologies represent a significant leap in the flexibility and quality of speech synthesis. From prosody and emotion to speaker identity and style, these systems empower diverse applications in fields such as entertainment, accessibility, and human-computer interaction.
@@
LLM-based synthesis. Here we pay special attention to LLM-based synthesis methods due to their superior context modeling capabilities compared to other neural-based TTS methods. LLMs, such as GPT [97, 98], T5 [99], and PaLM [100], have revolutionized various natural language processing (NLP) tasks with their ability to generate coherent, context-aware text. Recently, their utility has expanded into controllable TTS technologies [101, 102, 103, 104, 17]. For example, users can synthesize the target speech by describing its characteristics, such as: “A young girl says ‘I really like it, thank you!’ with a happy voice”, making speech generation significantly more intuitive and user-friendly. Specifically, an LLM can detect emotional intent in sentences (e.g., “I’m thrilled” → happiness, “This is unfortunate” → sadness). The detected emotion is encoded as an auxiliary input to the TTS model, enabling modulation of acoustic features like prosody, pitch, and energy to align with the expressed sentiment. By leveraging LLMs’ capabilities in understanding and generating rich contextual information, these systems can achieve enhanced and fine-grained control over various speech attributes such as prosody, emotion, style, and speaker characteristics [105, 31, 106]. Integrating LLMs into TTS systems represents a significant step forward, enabling more dynamic and expressive speech synthesis.
@@ I-C Organization of This Survey
This paper first presents a comprehensive and systematic review of controllable TTS technologies, with a particular focus on model architectures, control methodologies, and feature representations. To establish a foundational understanding, this survey begins with an introduction to the TTS pipeline in Section II. While our focus remains on controllable TTS, Section III examines seminal works in uncontrollable TTS that have significantly influenced the field’s development. Section IV provides a thorough investigation into controllable TTS methods, analyzing both their model architectures and control strategies. Section V presents a comprehensive review of datasets and evaluation metrics. Section VI provides an in-depth analysis of the challenges encountered in achieving controllable TTS systems and discusses future directions. Section VII explores the broader impacts of controllable TTS technologies and identifies promising future research directions, followed by the conclusion in Section VIII.
@@ II TTS Pipeline
Figure 2: The general pipeline of controllable TTS from the perspective of network structure. Linguistic analysis is necessary for parametric and a few neural-based methods but is no longer needed for most modern neural-based methods. In this paper, we only review neural-based controllable TTS methods and do not investigate acoustic features (e.g., MCC [107], LSP [108], F0 [109]) used in early TTS methods.
In this section, we elaborate on the general pipeline that supports controllable TTS technologies, including acoustic models, speech vocoders, and feature representations. Fig.2 depicts the general pipeline of controllable TTS, containing various model architectures and feature representations, but the control strategies will be discussed in Section IV. Readers can jump to Section III if familiar with TTS pipelines.
@@ II-A Overview
A TTS pipeline generally contains three key components, i.e., linguistic analyzer, acoustic model, speech vocoder, and with a conditional input, e.g., prompts, for controllable speech synthesis. Besides, some end-to-end methods use a single model to encode the input and decode the speech waveforms without generating intermediate features like mel-spectrograms [110]. Linguistic analyzer aims to extract linguistic features, e.g., phoneme duration and position, syllable stress, and utterance level, from the input text, which is a necessary step in HHM-based methods [64, 65] and a few neural-based methods [111, 112], but is time-consuming and error-prone. Acoustic model is a parametric or neural model that predicts the acoustic features from the input texts. Modern neural-based acoustic models like Tacotron [74] and later works [15, 76, 113] directly take character [114] or word embeddings [115] as the input, which is much more efficient than previous methods. Speech vocoder is the last component that converts the intermediate acoustic features into a waveform that can be played back. This step bridges the gap between the acoustic features and the actual sounds produced, helping to generate high-quality, natural-sounding speech [73, 116]. Tan et al. [42] have presented a comprehensive and detailed review of acoustic models and vocoders. Therefore, the following subsections will briefly introduce some representative acoustic models and speech vocoders, followed by a discussion of acoustic feature representations.
@@ II-B Acoustic Models
Acoustic modeling is a crucial step in TTS because it ensures the generated acoustic features capture the subtleties of human speech. By accurately modeling acoustic features, modern TTS systems can help generate high-quality and expressive audio that sounds close to human speech.
@@
RNN-based models. Recurrent Neural Networks (RNNs) proved particularly effective in early neural-based TTS due to their ability to model sequential data and long-range dependencies, which helps in capturing the sequential nature of speech, such as the duration and natural flow of phonemes. Typically, these models have an encoder-decoder architecture, where an encoder encodes input linguistic features, such as phonemes or text, into a fixed-dimensional representation, and the decoder sequentially decodes this representation into acoustic features (e.g., mel-spectrogram frames) that capture the frequency and amplitude of sound over time. Tacotron 2 [75] is one of the pioneering TTS models that uses RNNs with an attention mechanism, which helps align the text sequence with the generated acoustic features. It takes raw characters as input and produces mel-spectrogram frames, which are subsequently converted to waveforms. Another example is MelNet [121], which leverages autoregressive modeling to generate high-quality mel-spectrograms, demonstrating versatility in generating both speech and music, achieving high fidelity and coherence across temporal scales.
@@
CNN-based models. Unlike RNNs, which process sequential data frame by frame, CNNs process the entire sequence at once by applying filters across the input texts. This parallel approach enables faster training and inference, making CNN-based TTS particularly appealing for real-time and low-latency applications. Furthermore, by stacking multiple convolutional layers with varying kernel sizes or dilation rates, CNNs can capture both short-range and long-range dependencies, which are essential for natural-sounding speech synthesis. Deep Voice [122] is one of the first prominent CNN-based TTS models by Baidu, designed to generate mel-spectrograms directly from phoneme or character input. ParaNet [123] also utilizes a RNN model to achieve sequence-to-sequence mel-spectrogram generation. It uses a non-autoregressive architecture, which enables significantly faster inference by predicting multiple time steps simultaneously.
@@
Transformer-based models. Transformer model [124] uses self-attention layers to capture relationships within the input sequence, making them well-suited for tasks requiring an understanding of global contexts, such as prosody and rhythm in TTS. Transformer-based TTS models often employ an encoder-decoder architecture, where the encoder processes linguistic information (e.g., phonemes or text) and captures contextual relationships, and the decoder generates acoustic features (like mel-spectrograms) from these encoded representations, later converted to waveforms by a vocoder. TransformerTTS [125] is one of the first TTS models that apply transformers to synthesize speech from text. It utilizes a standard encoder-decoder transformer architecture and relies on multi-head self-attention mechanisms to model long-term dependencies, which helps maintain consistency and natural flow in speech over long utterances. FastSpeech [15] is a non-autoregressive model designed to overcome the limitations of autoregressive transformers in TTS, achieving faster synthesis than previous methods. It introduces a length regulator to align text with output frames, enabling the control of phoneme duration. FastSpeech 2 [76] extends FastSpeech by adding pitch, duration, and energy predictors, resulting in more expressive and natural-sounding speech.
@@
Other acoustic models. In TTS, GANs [127, 128, 129], VAEs [130, 18], and diffusion models [113, 131] can also be used as acoustic models. Flow-based methods [132, 133] are also popular in waveform generation. Refer to the survey paper from Tan et al. [42] for more details.

### v2

@@ Abstract
Text-to-speech (TTS), also known as speech synthesis, is a prominent research area that aims to generate natural-sounding human speech from text. Recently, with the increasing industrial demand, TTS technologies have evolved beyond synthesizing human-like speech to enabling controllable speech generation. This includes fine-grained control over various attributes of synthesized speech such as emotion, prosody, timbre, and duration. In addition, advancements in deep learning, such as diffusion and large language models, have significantly enhanced controllable TTS over the past several years. In this work, we conduct a comprehensive survey of controllable TTS, covering approaches ranging from basic control techniques to methods utilizing natural language prompts, aiming to provide a clear understanding of the current state of research. We examine the general controllable TTS pipeline, challenges, model architectures, and control strategies, offering a comprehensive and clear taxonomy of existing methods. Additionally, we provide a detailed summary of datasets and evaluation metrics and shed some light on the applications and future directions of controllable TTS. To the best of our knowledge, this survey paper provides the first comprehensive review of emerging controllable TTS methods, which can serve as a beneficial resource for both academic researchers and industrial practitioners.
@@ I Introduction
Deep learning [12] has made great progress in the past decade due to exponentially growing computational resources like GPUs [13], leading to the explosion of numerous exciting works on TTS [14, 15, 16, 17]. These methods can synthesize human speech with improved quality [14] and can achieve fine-grained control of the generated voice [18, 19, 20, 21, 22]. In addition, some recent works synthesize speech given multi-modal input, such as face images [23, 24], cartoons [7], and videos [25]. Moreover, with the fast development of open-source LLMs [11, 26, 27, 28, 29], some researchers propose to synthesize fine-grained controllable speech with natural language description [30, 31, 32], offering a new way to generate custom speech voices. Meanwhile, powering LLMs with speech synthesis has also been a hot topic in the last few years [33, 34, 35]. In recent years, a wide range of TTS methods has emerged, making it essential for researchers to gain a comprehensive understanding of current research trends, particularly in controllable TTS, and to identify promising future directions in this rapidly evolving field. Consequently, there is a pressing need for an up-to-date survey of TTS techniques. While several existing surveys address parametric approaches [36, 37, 38, 39, 40, 41] and deep learning-based approaches [42, 43, 44, 45, 46, 47, 48], they largely overlook the controllability of TTS. Additionally, these surveys do not cover recent advancements, such as natural language description-based TTS methods.
This paper provides a comprehensive and in-depth survey of existing and emerging TTS technologies, with a particular focus on controllable TTS methods. Fig. 1 demonstrates the development of controllable TTS methods in recent years, showing their backbones, feature representations, and control abilities. The remainder of this section begins with a brief comparison between this survey and previous ones, followed by an overview of the history of controllable TTS technologies, ranging from early milestones to state-of-the-art advancements. Finally, we introduce the taxonomy and organization of this paper. We have posted a version of our paper on arXiv.org (https://arxiv.org/abs/2412.06602).
Figure 1: A summary of representative controllable TTS methods in recent years and their model architectures, feature representations, and control abilities. Additional network structures, such as VAE and flow-based models, are not included in this figure. For more details, refer to Tables IV and III.
@@ I-A Comparison with Existing Surveys
Different Scope. Klatt et al. [36] provided the first comprehensive survey on formant, concatenative, and articulatory TTS methods, with a strong emphasis on text analysis. In the early 2010s, Tabet et al. [49] and King et al. [40] explored rule-based, concatenative, and Hidden Markov Models (HMM)-based techniques. Later, the advent of deep learning catalyzed the emergence of numerous neural model-based TTS methods. Therefore, Ning et al. [43] and Tan et al. [42] have conducted extensive surveys on neural acoustic models and vocoders, while Zhang et al. [50] presented the first review of diffusion model-based TTS techniques. However, these studies offer limited discussion on the controllability of TTS systems. To address this gap, we present the first comprehensive survey of TTS methods through the lens of controllability, providing an in-depth analysis of model architectures and strategies for controlling synthesized speech.
Close to Current Demands. With the rapid development of hardware (i.e., GPUs and TPUs) and artificial intelligence (AI) techniques (i.e., transformers, LLMs, diffusion models) in the last few years, the demand for controllable TTS is becoming increasingly urgent due to its broad applications in industries such as filmmaking, gaming, robotics, and personal assistants. Despite this growing need, existing surveys pay little attention to control methods in TTS technologies. To bridge this gap, we propose a systematic analysis of current controllable TTS methods and the associated challenges, offering a comprehensive understanding of the research state in this field.
@@
Table I summarizes the existing surveys and our survey in terms of main focus and publication year.
@@
Tan et al. [42] Neural, efficient, and expressive TTS 2021
@@
Ours Controllable TTS, datasets, metrics, and challenges 2025
@@ I-B The History of Controllable TTS
Controllable TTS aims to control various aspects of synthesized speech, such as pitch, energy, speed/duration, prosody, timbre, emotion, gender, or high-level styles. This subsection briefly reviews the history of controllable TTS, ranging from early approaches to the state-of-the-art (SOTA) in recent years.
@@
Early Approaches. Before the prevalence of deep neural networks (DNNs), controllable TTS technologies were built primarily on rule-based, concatenative, and statistical methods. These approaches enable some degree of customization and control, however, they were constrained by the limitations of the underlying models and available computational resources. 1) Rule-based TTS systems [51, 52, 53, 54], such as formant synthesis, were among the earliest methods for speech generation. These systems use manually crafted rules to simulate the speech generation process by controlling acoustic parameters such as pitch, duration, and formant frequencies, allowing explicit manipulation of prosody and phonetic details through rule adjustments. 2) Concatenative TTS [55, 56, 57, 58], which dominated the field in the late 1990s and early 2000s, synthesize speech by concatenating pre-recorded speech segments, such as phonemes or diphones, stored in a large database [59]. These methods can modify the prosody by manipulating the pitch, duration, and amplitude of speech segments during concatenation. They also allow limited voice customization by selecting speech units from different speakers. 3) Parametric methods, particularly HMM-based TTS [60, 61, 62, 63, 64, 65], gained prominence in the late 2000s. These systems model the relationships between linguistic features and acoustic parameters, providing more flexibility in controlling prosody, pitch, speaking rate, and timbre by adjusting statistical parameters. Some HMM-based systems also supported speaker adaptation [66, 67] and voice conversion [68, 69], enabling voice cloning to some extent. However, emotion can be limitedly controlled by some of these methods [70, 71, 60, 72]. In addition, they required less storage compared to concatenative TTS and allowed smoother transitions between speech units.
@@
Neural Synthesis. Neural model-based TTS technologies emerged with the advent of deep learning, significantly advancing the field by enabling more flexible, natural, and expressive speech synthesis. Unlike traditional methods, neural TTS leverages DNNs to model complex relationships between input text and speech, facilitating nuanced control over various speech characteristics. Early neural TTS systems, such as WaveNet [73] and Tacotron [74], laid the groundwork for controllability. 1) Controlling prosody features like rhythm and intonation is vital for generating expressive and contextually appropriate speech. Neural TTS models achieve prosody control through explicit conditioning or learned latent representations [75, 15, 76, 77, 78]. 2) Speaker control has also gained significant improvement in neural TTS through speaker embeddings or adaptation techniques [79, 80, 81, 82]. 3) Besides, emotionally controllable TTS [83, 31, 20, 22, 32] has become a hot topic due to the strong modeling capability of DNNs, enabling the synthesis of speech with specific emotional tones, such as happiness, sadness, anger, or neutrality. These systems go beyond producing intelligible and natural-sounding speech, focusing on generating expressive output that aligns with the intended emotional context. 4) Neural TTS can also manipulate timbre (vocal quality) [78, 84, 85, 14, 86, 87] and style (speech mannerisms) [88, 89, 90], allowing for creative and personalized applications. These techniques lead to one of the most popular research topics, i.e., zero-shot TTS (particularly voice cloning) [82, 78, 91, 92]. 5) Fine-grained content and linguistic control also become more powerful [93, 94, 95, 96]. These methods can emphasize or de-emphasize specific words or adjust the pronunciation of phonemes through speech editing or generation techniques.
@@
Neural TTS technologies represent a significant leap in the flexibility and quality of speech synthesis. From prosody and emotion to speaker identity and style, these systems empower diverse applications in fields such as entertainment, accessibility, and human-computer interaction.
@@
LLM-based Synthesis. Here, we pay special attention to LLM-based synthesis methods due to their superior context modeling capabilities compared to other neural TTS methods. LLMs, such as generative pre-trained transformer (GPT) [97, 98], T5 [99], and pathways language model (PaLM) [100], have revolutionized various natural language processing (NLP) tasks with their ability to generate coherent, context-aware text. Recently, their utility has expanded into controllable TTS technologies [101, 102, 103, 104, 17]. For example, users can synthesize the target speech by describing its characteristics, such as: “A young girl says ‘I really like it, thank you!’ with a happy voice”, making speech generation significantly more intuitive and user-friendly. Specifically, an LLM can detect emotional intent in sentences (e.g., “I’m thrilled” → happiness, “This is unfortunate” → sadness). The detected emotion is encoded as an auxiliary input to the TTS model, enabling the modulation of acoustic features like prosody, pitch, and energy to align with the expressed sentiment. By leveraging LLMs’ capabilities in understanding and generating rich contextual information, these systems can achieve enhanced and fine-grained control over various speech attributes, such as prosody, emotion, style, and speaker characteristics [105, 31, 106]. Integrating LLMs into TTS represents a significant step forward, enabling more dynamic and expressive speech synthesis.
@@ I-C Organization of This Survey
This survey first presents a comprehensive and systematic review of controllable TTS technologies, with a particular focus on model architectures, control strategies, and feature representations. To establish a foundational understanding, this survey begins with an introduction to the TTS pipeline in Section II. While our focus remains on controllable TTS, Section III examines seminal works in uncontrollable TTS that have significantly influenced the field’s development. Section IV provides a thorough investigation into controllable TTS methods, analyzing both their model architectures and control strategies. Section V presents a comprehensive review of datasets and evaluation metrics. Section VI provides an in-depth analysis of the challenges encountered in achieving controllable TTS systems and identifies promising future research directions. Section VII explores the broader impacts of controllable TTS technologies, followed by the conclusion in Section VIII.
@@ II TTS Pipeline
Figure 2: General pipeline of controllable TTS from the perspective of network structure. Linguistic analysis is necessary for parametric and a few neural methods but is no longer needed for most modern neural methods. In this paper, we only review neural model-based controllable TTS methods and do not investigate acoustic features (e.g., MFCC [107], LSP [108], F0 [109]) used in early TTS methods.
In this section, we elaborate on the general pipeline that supports controllable TTS technologies, including acoustic models, speech vocoders, and feature representations. Fig. 2 depicts the general pipeline of controllable TTS, containing various model architectures and feature representations, but the control strategies will be discussed in Section IV. Readers familiar with TTS pipelines may skip ahead to Section III.
@@ II-A Overview
A TTS pipeline generally contains three key components, i.e., linguistic analyzer, acoustic model, and speech vocoder, where a conditional input, e.g., prompts, can be processed for controllable speech synthesis. Linguistic analyzer aims to extract linguistic features, e.g., phoneme duration and position, syllable stress, and utterance level, from the input text, which is a necessary step in HMM-based methods [64, 65] and a few neural model-based methods [110, 111], but is time-consuming and error-prone. Acoustic model is a parametric or neural model that predicts the acoustic features from the input texts. Modern neural acoustic models like Tacotron [74] and later works [15, 76, 112] directly take character [113] or word embeddings [114] as the input, which is much more efficient than previous methods. Speech vocoder is the last component that converts the intermediate acoustic features into a waveform that can be played back. This step bridges the gap between the acoustic features and the actual sounds produced, helping to generate high-quality, natural-sounding speech [73, 115]. Besides, some end-to-end methods use a single model to encode the input and decode the speech waveforms without generating intermediate features like mel-spectrograms. Tan et al. [42] have presented a comprehensive and detailed review of acoustic models and vocoders. Therefore, the following subsections will briefly introduce some representative acoustic models and speech vocoders, followed by a discussion of acoustic feature representations.
@@ II-B Acoustic Models
Acoustic modeling is a crucial step in TTS because it ensures that the generated acoustic features capture the subtleties of human speech. By accurately modeling acoustic features, modern TTS systems can help generate high-quality and expressive audio that sounds close to human speech.
@@
RNN-based Models. Recurrent Neural Networks (RNNs) proved particularly effective in early neural TTS due to their ability to model sequential data and long-range dependencies, which helps in capturing the sequential nature of speech, such as the duration and natural flow of phonemes. Typically, these models have an encoder-decoder architecture, where an encoder encodes input linguistic features, such as phonemes or text, into a fixed-dimensional representation, and the decoder sequentially decodes this representation into acoustic features (e.g., mel-spectrogram frames) that capture the frequency and amplitude of sound over time. Tacotron 2 [75] is one of the pioneering TTS models that use RNNs with an attention mechanism, which helps align the text sequence with the generated acoustic features. It takes raw characters as input and produces mel-spectrogram frames, which are subsequently converted to waveforms. Another example is MelNet [119], which leverages autoregressive modeling to generate high-quality mel-spectrograms, demonstrating versatility in generating both speech and music, achieving high fidelity and coherence across temporal scales.
@@
CNN-based Models. Unlike RNNs, which process sequential data frame by frame, CNNs process the entire sequence at once by applying filters across the input texts. This parallel approach enables faster training and inference, making CNN-based TTS particularly appealing for real-time and low-latency applications [73, 16, 120, 121]. Furthermore, by stacking multiple convolutional layers with varying kernel sizes or dilation rates, CNNs can capture both short-range and long-range dependencies, which are essential for natural-sounding speech synthesis. Deep Voice [16] is one of the first prominent CNN-based TTS methods, designed to generate mel-spectrograms directly from phoneme or character input. ParaNet [122] also utilizes a CNN model to achieve sequence-to-sequence mel-spectrogram generation. It uses a non-autoregressive architecture, which enables significantly faster inference by predicting multiple time steps simultaneously.
@@
Transformer-based Models. The transformer model [123] uses self-attention layers to capture relationships within the input sequence, making them well-suited for tasks requiring an understanding of global contexts, such as prosody and rhythm in TTS. Transformer-based TTS models often employ an encoder-decoder architecture, where the encoder processes linguistic information (e.g., phonemes or text) and captures contextual relationships, and the decoder generates acoustic features (like mel-spectrograms) from these encoded representations, later converted to waveforms by a vocoder. TransformerTTS [124] is one of the first TTS models that apply transformers to synthesize speech from text. It utilizes a standard encoder-decoder transformer architecture and relies on multi-head self-attention mechanisms to model long-term dependencies, which helps maintain consistency and natural flow in speech over long utterances. FastSpeech [15] is a non-autoregressive model designed to overcome the limitations of autoregressive transformers in TTS, achieving faster synthesis than previous methods. It introduces a length regulator to align text with output frames, enabling the control of phoneme duration. FastSpeech 2 [76] extends FastSpeech by adding pitch, duration, and energy predictors, resulting in more expressive and natural-sounding speech.
@@
Other Acoustic Models. In TTS, generative adversarial networks (GANs) [126, 127, 128], variational autoencoders (VAEs) [129, 18], and diffusion models [112, 130] can also be used as acoustic models. Flow-based methods [131, 132] are also popular in acoustic feature generation. Recently proposed flow-based controllable TTS methods will be discussed in Section IV. Refer to the survey paper from Tan et al. [42] for more details.